A 100 ohm resistor is connected in series in a parallel combination of three parallel resistors 4 ohm 10 ohm and 20 ohm respectively and its connected to a 50 volt DC supply. Find the current in each resistor and total power dissipated.
1
Expert's answer
2020-11-20T09:35:20-0500

As per question ,

Let the Resistors be,


"R_1=100\Omega,R_2=4\Omega,R_3=10\Omega" and "R_4=20\Omega"


As per question,

"R_1" is connected in series with parallel combination of "R_2,R_3,R_4"


Net resistance "R_{net}=R_1+\dfrac{1}{\dfrac{1}{R_2}+\dfrac{1}{R_2}+\dfrac{1}{R_3}}"


"\Rightarrow R_{net}=100+\dfrac{1}{\dfrac{1}{4}+\dfrac{1}{10}+\dfrac{1}{20}}"


"\Rightarrow R_{net}=100+\dfrac{1}{\dfrac{5+2+1}{20}}"


"\Rightarrow R_{net}=100+\dfrac{20}{8}"


"\Rightarrow R_{net}=100+2.5=102.5\Omega"


Current in the circuit "i=\dfrac{V}{R_{net}}" ( Ohm's law)


"i=\dfrac{50}{102.5}=0.4878A"


Current in resistor "R_1" is "i_1=i=0.4878A"

Power dissipation "P_1=i_1^2R_1=(0.4878)^2(100)=23.79 watt"


Voltage drops across "R_1" resistor "V_1=i\times R_1=0.4878\times 100=48.78V"


Now the remaining voltage "V_2=V-V_1=50-48.78=1.22V"


Since resistor "R_2,R_3,R_4" are in parallel so Voltage remains same,


"\therefore" Current in resistor "R_2" is "i_2=\dfrac{1.22}{4}=0.305" A

Power dissipation "P_2=i_2^2R_2=(0.305)^2(4)=0.3721 watt"


"\therefore" Current in resistor "R_3" is "i_3=\dfrac{1.22}{10}=0.122A"

Power dissipation "P_3=i_3^2R_3=(0.122)^2(10)=0.14 watt"


"\therefore" Current in resistor "R_4" is "i_4=\dfrac{1.22}{20}=0.061A"

Power dissipation "P_4=i_4^2R_4=(0.061)^2(20)= 0.07442 watt"